OpenStudy (anonymous):

what is the probability that at least 2 will survive if expectancy for 3 people are 0.1,0.2,0.3

OpenStudy (anonymous):

1st we need to clear up what those probabilities represent. You said "expectancy".Are those their probabilities of dying or living?

OpenStudy (anonymous):

living

OpenStudy (anonymous):

PR(AdiesBlivesClives) + Pr(AlivesBdiesClives) + Pr(AlivesBlivesCdies + Pr(all live) 0.9*0.2*0.3 + 0.1*0.8*0.3 + 0.1*0.2*0.7 + 0.1*0.2*0.3

OpenStudy (anonymous):

For each person, the probability they die is 1 - Pr(live) 1-0.1=0.9 1-0.2=0.8 1-0.3=0.7

OpenStudy (anonymous):

okay so we use the compliment correct?

OpenStudy (anonymous):

Yes...for the probability of dying.

OpenStudy (anonymous):

i got 0.098 for the total. is that the answer or do i have to do something else?

OpenStudy (anonymous):

That's what I get, too

OpenStudy (anonymous):

These people must have some terrible disease to have such small chances of living!

OpenStudy (anonymous):

i know right haha these are just the numbers my teacher gave me

OpenStudy (anonymous):

so how would i find the probability that only 2 will survive

OpenStudy (anonymous):

Same way...just leave off that last 0.1*0.2*0.3...which was the prob that all 3 live.

OpenStudy (anonymous):

alright thank you!

OpenStudy (anonymous):

No problem.

OpenStudy (anonymous):

A little secret. I actually used to work for a life insurance where we actually had to consider the probabilities of people dying when setting life insurance rates. So these types of questions are VERY familiar to me

OpenStudy (anonymous):

did you get 0.092?

OpenStudy (anonymous):

oh really?! hopefully you werent getting numbers like these were you?

OpenStudy (anonymous):

oops! yes you are correct.

OpenStudy (anonymous):

oh alright i just rechecked it all i was confused haha